# Panagiotis Koletsis Dept. of Informatics and Telematics

### v1

@@ I Introduction
Financial crimes have always been a concern, but the recent geopolitical instabilities have played a crucial role in a significant increase of illicit financial activities, including sanctions evasion, money laundering, and the use of shadowy financial networks to fund conflicts and circumvent international restrictions [1]. Geopolitical tensions have created opportunities for criminals to exploit the chaos and for sanctioned individuals and entities to find new ways to access global markets [2]. In general, most research work in financial data is focused on areas related to the stock market, such as market analysis [3] and stock price prediction [4]. The primary goal of this paper is to extract information related to financial crime data, and contribute to the limited literature in the topic. The information extraction is conducted on news articles that mentioned individuals and organizations. This information can then be used by experts in the investigation of financial crime cases.
The latest advances in Large Language Models (LLMs) have sparked a new family of algorithms that rely on prompt engineering to perform natural-language processing (NLP) tasks, like named entity recognition (NER) [5] or attribute value extraction [6]. Therefore, a natural question that arises is how well would an LLM with appropriate prompting work in the tasks described above. Due to the lack of research in this specific field, there are no datasets specialized in financial crime that can be used to evaluate the performance of the LLM responses. In this case, the interest is specifically focused on a list of named entities, individuals (persons) or organizations, mentioned within some provided articles. To address this, an automated pipeline for dataset generation is created without compromising reliability. The resulting dataset contains ground truth for individual and organization identification in JSON format for each included article. In addition, a helper LLM is used to fix the format of malformed JSON responses. Lastly, the issue of ambiguous entity mentions, (i.e., references to the same element in different ways) is tackled by developing a new pipeline that uses an LLM to extract matching elements.
@@
The source code of this work is publicly available on GitHub: https://github.com/panagiotis-koletsis/KleptotraceDataset.
@@ II Related Work
Financial Market Tasks. Most research regarding financial data is mainly focused in the area of financial markets, due to the wide range of interest that exists. This is true for both classical Machine Learning Methods and Deep Learning Methods [13]. Especially for Large Language Models the research in this area is currently very active. Starting with FinBERT [14], a transformer model based on BERT, which primarily focuses on sentiment analysis of financial data, achieving state-of-the-art performance on related datasets, such as the Financial Phrase Bank in 2019 [15, 16]. In 2023, Bloomberg launched its own LLM based on the BLOOM architecture, called BloombergGPT, an enterprise AI assistant for financial tasks such as investment decisions and risk management [17]. Shortly after, the need for an open-source version of BloombergGPT is met with the release of FinGPT, providing similar capabilities to the Bloomberg counterpart [18].
General Tasks. Named Entity Recognition (NER) has been a trending subject for the past few decades [19]. Initially, statistical models provided acceptable results [20]. More recently, a long short-term memory (LSTM)-based model called Flair embeddings [21] achieved a 93.09 F1 score on the CoNLL-2003 Hugging Face dataset [^1] https://huggingface.co/datasets/eriktks/conll2003. With the introduction of the attention layer and the implementation of transformer models, results are further improved [22]. A BERT-based model called LUKE achieved a 94.3 F1 score on the same dataset [23]. The state-of-the-art performance on this dataset is currently achieved not by a single model, but by a technique that uses reinforcement learning to search for the best embedding among multiple candidates provided by different models, named ACE [24].
@@ III-A Micro-Benchmark Dataset Generation
The benchmark generation methodology receives as input a set of articles and returns annotations, as lists for those articles, reflecting the named entities (individuals and organizations). In this use case, the benchmark generation pipeline takes as input a large dataset containing multiple articles from various cases, that is provided within the context of EU KLEPTOTRACE project [25]. An example is shown in (Figure 2), highlighting the annotations (detected individuals and organizations) returned for a given article. For presentation purposes, only a snippet of this article is shown in the figure.
@@ III-A1 Annotation Process
  • spaCy annotation: The spaCy library is utilized to assist with manual annotation
  • Manual annotation
  • Final verification by a Large Language Model (LLM) using ChatGPT 3.5.
  • To ensure accuracy in identifying the organizations, answers are validated through comprehensive web searches when needed
@@ III-A2 Data Statistics and Description
The generated dataset contains 15 randomly selected articles from the provided set of articles, annotated in a JSON format. The articles vary in length from 170 to 1,467 words. Each article contains a different number of Individuals/Organizations to be identified, ranging from 1 to 12 for Individuals and 0 to 16 for Organizations. In total, there are 84 individuals and 128 organizations mentioned in the dataset. Each article has different characteristics, allowing the proposed methodology to be tested under various circumstances. The final form of the dataset consists of 441 sentences, 11,152 words, and 72,332 characters.
@@ III-B Evaluation Methodology
The LLM-based extraction used for evaluation purposes, as well as the evaluation process, are shown in (Figure 4) and described in detail, in the following.
@@
Next, the matching elements from List 1 and List 2 are renamed in List 2, if needed, to use the same names as those used in List 1. (see Renaming in Figure 1). Finally, the updated LLM List is taken, which contains the initial elements from LLM List, with the common elements renamed as they appear in the Ground Truth list.
@@ IV-A Experimental Setting
Dataset: The dataset used is the one referenced in section [III.A], which contains articles related to financial crimes and all the mentioned individuals and organizations. It will be made publicly available after the paper acceptance.
@@ IV-B Experimental Results
In this section, we report the evaluation results on the NER task using the LLM prompts described in Section III-B, on the new benchmark dataset described in Section III-A. More specifically, in Tables I and II, we present the results of Individual Identification and Organization identification, respectively, with the last column showing the additions made to the base prompt (shown in Figures 10 and  11). In Table III we compare those results to a baseline method (spaCy) and in Table IV we discussing the results of the Structuring LLM component.
@@
qwen2:7b 0.342 0.764 0.382 0.508 74.9 -
llama3:8b 0.514 0.851 0.565 0.678 83.5 -
gemma:7b 0.523 0.767 0.622 0.687 114.5 -
gemma2:9b 0.640 0.961 0.657 0.780 36.2 -
gemma2:9b 0.663 0.941 0.692 0.797 37.6 1)
gemma2:9b 0.636 0.914 0.676 0.777 37.2 2)
gemma2:9b 0.612 0.913 0.650 0.759 36.5 3)
gemma2:9b 0.683 0.949 0.709 0.811 48.8 4)
gemma2:9b 0.653 0.946 0.679 0.790 37.9 5)
gemma2:9b 0.699 0.936 0.734 0.823 59.5 1),4),5)
TABLE III: Comparison with spaCy Framework.
Base model Accuracy Precision Recall F1 Score Execution Time (Sec) Experiment Matching
spaCy 0.447 0.486 0.845 0.617 0.9 Individual -
spaCy 0.117 0.176 0.258 0.209 0.9 Organization -
spaCy 0.297 0.393 0.549 0.458 27.7 Organization yes
TABLE IV: Individual identification for Structuring LLM.
@@
  • The best precision was achieved with the zero-shot (base prompt) approach, but it also resulted in one of the worst recalls. Additionally, it had the lowest iteration time.
@@
SpaCy comparison: The proposed methodology was compared with spaCy (Table III), as it was employed directly in the dataset creation process. Spacy is a framework commonly used in industrial NLP processes. The en_core_web_sm model from spaCy was specifically used and tested on the developed dataset to compare results. First, the spaCy framework with the raw model outputs was evaluated, and as expected, the results from the new proposed method were superior. Organization identification was also tested with the spaCy framework, including the matching component, and found that the suggested approach still outperformed spaCy. One notable advantage of using spaCy is its rapid output. In less than one second, spaCy was able to process all 15 of articles.
Structuring results: To evaluate the Structuring LLM, the entire individual recognition experiment was ran 5 times, counting JSON errors and failure percetange, as shown in Table IV. This equates to 5 x 7 x 15 = 525 iterations, because each experiment involved 7 iterations of parsing each article, and there were 15 articles in total. Gemma2:9b had 4 JSON errors, indicating that 4 out of the 525 responses were not structured properly. However, in the conducted experiments, Qwen2:7B achieved perfect structuring with 0 JSON errors.
Discussion: Overall, one interesting finding was that both Gemma2:27B and Phi3 (which has 14 billion parameters), despite being recent and powerfull models in other NLP tasks, seemed to struggle with these tasks, as much as other, older models. Another observation was that, despite the initial unsatisfactory results of Qwen2:7B, zero JSON errors were encountered consistently with this model. This indicates that Qwen2:7B understood the structure of the format better than other models, such as LLaMA3:8B and Gemma:7B, which struggled significantly more. Therefore, Qwen2:7B was used as the Structuring LLM. However, Gemma2:9B outperformed other models (LLaMA3:8B, Gemma:7B) in structuring,which was empirically proven. Nevertheless, it did not match the performance of Qwen2:7B, which excelled at structuring its responses, as demonstrated experimentally (Table IV).

### v2

@@ I Introduction
Financial crimes have always been a concern, but the recent geopolitical instabilities have played a crucial role in a significant increase of illicit financial activities, including sanctions evasion, money laundering, and the use of shadowy financial networks to fund conflicts and circumvent international restrictions [1]. Geopolitical tensions have created opportunities for criminals to exploit the chaos and for sanctioned individuals and entities to find new ways to access global markets [2]. In general, most research work in financial data is focused on areas related to the stock market, such as market analysis [3] and stock price prediction [4]. The primary goal of this paper is to extract information related to financial crime data, and contribute to the limited literature in the topic. The information extraction is conducted on news articles that mention individuals and organizations. This information can then be used by experts in the investigation of financial crime cases.
The latest advances in LLMs have sparked a new family of algorithms that rely on prompt engineering to perform natural-language processing (NLP) tasks, like named entity recognition (NER) [5] or attribute value extraction [6]. Therefore, a natural question that arises is how well would an LLM with appropriate prompting work in the tasks described above. Due to the lack of research in this specific field, there are no datasets specialized in financial crime that can be used to evaluate the performance of the LLM responses. In this case, the interest is specifically focused on a list of named entities, individuals (persons) or organizations, mentioned within some provided articles. To address this, an innovative pipeline for dataset generation has been created without compromising reliability. The resulting dataset contains ground truth for individual and organization identification in JSON format for each included article. In addition, a helper LLM is used to fix the format of malformed JSON responses. Lastly, the issue of ambiguous entity mentions, (i.e., references to the same element in different ways) is tackled by developing a new pipeline that uses an LLM to extract matching elements.
@@
The source code of this work is publicly available on GitHub: https://github.com/panagiotis-koletsis/KleptotraceDataset
@@ II Related Work
Financial Market Tasks. Most research regarding financial data is mainly focused in the area of financial markets, due to the wide range of interest that exists. This is true for both classical Machine Learning Methods and Deep Learning Methods [13]. Especially for LLMs the research in this area is currently very active. Starting with FinBERT [14], a transformer model based on BERT, which primarily focuses on sentiment analysis of financial data, achieving state-of-the-art performance on related datasets, such as the Financial Phrase Bank in 2019 [15, 16]. In 2023, Bloomberg launched its own LLM based on the BLOOM architecture, called BloombergGPT, an enterprise AI assistant for financial tasks such as investment decisions and risk management [17]. Shortly after, the need for an open-source version of BloombergGPT is met with the release of FinGPT, providing similar capabilities to the Bloomberg counterpart [18].
General Tasks in NER Datasets. NER has been a trending subject for the past few decades [19]. Initially, statistical models provided acceptable results [20]. More recently, a long short-term memory (LSTM)-based model called Flair embeddings [21] achieved a 93.09 F1 score on the CoNLL-2003 Hugging Face dataset [^1] https://huggingface.co/datasets/eriktks/conll2003. With the introduction of the attention layer and the implementation of transformer models, results are further improved [22]. A BERT-based model called LUKE achieved a 94.3 F1 score on the same dataset [23]. The state-of-the-art performance on this dataset is currently achieved not by a single model, but by a technique that uses reinforcement learning to search for the best embedding among multiple candidates provided by different models, named ACE [24].
@@ III-A Micro-Benchmark Dataset Generation
The benchmark generation methodology receives as input a set of articles and returns annotations, as lists for those articles, reflecting the named entities (individuals and organizations). In this use case, the benchmark generation pipeline takes as input a large dataset containing multiple articles from various cases, that is provided within the context of EU KLEPTOTRACE project [25]. This large unstructured dataset is comprised of more than 5000 multilingual articles from over 20 cases, collected from trusted sources, such as law enforcement press releases. An example is shown in (Figure 2), highlighting the annotations (detected individuals and organizations) returned for a given article. For presentation purposes, only a snippet of this article is shown in the figure.
@@ III-A1 Annotation Process
  • spaCy annotation: The spaCy library is utilized to assist with manual annotation.
  • Manual annotation: Inspect, correct and expand spaCy’s annotations.
  • Final verification by a LLM specifically ChatGPT 3.5.
  • To ensure accuracy in identifying the organizations, answers are validated through comprehensive web searches when needed.
@@ III-A2 Data Statistics and Description
The generated dataset contains 15 randomly selected articles from the provided set of articles, annotated in a JSON format.
The articles vary in length from 170 to 1,467 words. Each article contains a different number of Individuals/Organizations to be identified, ranging from 1 to 12 for Individuals and 0 to 16 for Organizations. In total, there are 84 individuals and 128 organizations mentioned in the dataset.
Each article has different characteristics, allowing the proposed methodology to be tested under various circumstances. The final form of the dataset consists of 441 sentences, 11,152 words, and 72,332 characters.
@@ III-B Evaluation Methodology
The LLM-based extraction process, as well as the evaluation process, are shown in Figure 4 and described in detail, below.
@@
Next, the matching elements from List 1 and List 2 are renamed in LLM list, if needed, to use the same names as those used in List 1. (see Renaming in Figure 1). Finally, the updated LLM List is taken, which contains the initial elements from LLM List, with the common elements renamed as they appear in the Ground Truth list.
@@ IV-A Experimental Setting
Dataset: The dataset used is the one referenced in section [III.A], which contains articles related to financial crimes and all the mentioned individuals and organizations.
@@ IV-B Experimental Results
In this section, we report the evaluation results on the NER task using the LLM prompts described in Section III-B, on the new benchmark dataset described in Section III-A. More specifically, in Tables I and II, we present the results of Individual identification and Organization identification, respectively, with the "Prompt Addition" column showing the additions made to the base prompt (shown in Figures 10 and  11). These tables also present a one-to-one comparison with the spaCy framework. Finally in Table III we are discussing the results of the Structuring LLM component.
@@
Base model Accuracy Precision Recall F1 Score Execution Time (Sec) Prompt Additions Matching
qwen2:7b 0.342 0.764 0.382 0.508 74.9 - yes
llama3:8b 0.514 0.851 0.565 0.678 83.5 - yes
gemma:7b 0.523 0.767 0.622 0.687 114.5 - yes
gemma2:9b 0.640 0.961 0.657 0.780 36.2 - yes
gemma2:9b 0.663 0.941 0.692 0.797 37.6 1) yes
gemma2:9b 0.636 0.914 0.676 0.777 37.2 2) yes
gemma2:9b 0.612 0.913 0.650 0.759 36.5 3) yes
gemma2:9b 0.683 0.949 0.709 0.811 48.8 4) yes
gemma2:9b 0.653 0.946 0.679 0.790 37.9 5) yes
gemma2:9b 0.699 0.936 0.734 0.823 59.5 1),4),5) yes
spaCy 0.117 0.176 0.258 0.209 0.9 - no
spaCy 0.297 0.393 0.549 0.458 27.7 - yes
TABLE III: Individual identification for Structuring LLM.
@@
  • The best precision was achieved with the zero-shot (base prompt) approach, but it also resulted in one of the worst recalls. Additionally, it had the lowest iteration time among the LLM approaches.
@@
SpaCy comparison: The proposed methodology was compared with spaCy (Tables I and II), as it was employed directly in the dataset creation process. Spacy is a framework commonly used in industrial NLP processes. The en_core_web_sm model from spaCy was specifically used and tested on the developed dataset to compare results. First, the spaCy framework with the raw model outputs was evaluated, and as expected, the results from the new proposed method were superior. Organization identification was also tested with the spaCy framework, including the matching component, and found that the suggested approach still outperformed spaCy. One notable advantage of using spaCy is its rapid output. In less than one second, spaCy was able to process all 15 of articles.
Structuring results: To evaluate the Structuring LLM, the entire individual recognition experiment was ran 5 times, counting JSON errors and failure percetange, as shown in Table III. This equates to 5 x 7 x 15 = 525 iterations, because each experiment involved 7 iterations of parsing each article, and there were 15 articles in total. Gemma2:9b had 4 JSON errors, indicating that 4 out of the 525 responses were not structured properly. However, in the conducted experiments, Qwen2:7B achieved perfect structuring with 0 JSON errors.
Discussion: Overall, one interesting finding was that both Gemma2:27B and Phi3 (which has 14 billion parameters), despite being recent and powerfull models in other NLP tasks, seemed to struggle with these tasks, as much as other, older models. Another observation was that, despite the initial unsatisfactory results of Qwen2:7B, zero JSON errors were encountered consistently with this model. This indicates that Qwen2:7B understood the structure of the format better than other models, such as LLaMA3:8B and Gemma:7B, which struggled significantly more. Therefore, Qwen2:7B was used as the Structuring LLM. However, Gemma2:9B outperformed other models (LLaMA3:8B, Gemma:7B) in structuring,which was empirically proven. Nevertheless, it did not match the performance of Qwen2:7B, which excelled at structuring its responses, as demonstrated experimentally (Table III).
Limitations: Most of the limitations identified are resulting from the utilization of LLMs.
  • Relatively slow and GPU intensive process.
  • Suffering from hallucinations, which can affect the results and their trustworthiness.
  • The results are non-deterministic, as the output of LLMs may vary across different iterations.
  • Depending on third parties libraries for the LLM utilization, such as Ollama and Langchain.